40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 25 20:44:11 2012, Jenne, Update, SUS, Tip tilts in progress 
    Reply  Fri Oct 26 10:44:25 2012, Jenne, Update, SUS, Tip tilts in progress 
       Reply  Fri Oct 26 13:08:14 2012, Jenne, Update, SUS, Tip tilts in progress 
Message ID: 7631     Entry time: Fri Oct 26 13:08:14 2012     In reply to: 7630
Author: Jenne 
Type: Update 
Category: SUS 
Subject: Tip tilts in progress 

Quote:

Quote:
 

Jamie is working on re-pitch aligning TT004 and TT005 (we already did 001), then we can re-install them in the vacuum system later this afternoon.

 The tip tilts have all been pitch-adjusted now, and they have all been put back onto the tables, with the same serial numbers in the same places as we took them out.  Jamie also re-leveled the BS table.

Raji and I will align things after I finish measuring the MC spot positions.

ELOG V3.1.3-